OpenStudy (anonymous):

Fifty-three plus four times b is as much as 21. Help!!! Fan+ Medal!

OpenStudy (anonymous):

is it 53+4b=21?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

sry i dont know this one

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

you have the correct equation set up

OpenStudy (anonymous):

@jim_thompson5910 do you know?

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

So I am correct?

jimthompson5910 (jim_thompson5910):

I'm guessing they want you to solve for b? or just set up the equation?

OpenStudy (anonymous):

just setup

OpenStudy (anonymous):

or basically put it in number form

jimthompson5910 (jim_thompson5910):

then you have the correct answer they want

OpenStudy (anonymous):

The sum of 5 times h and twice g is equal to 23

OpenStudy (anonymous):

5h+g^2=23

OpenStudy (anonymous):

????

jimthompson5910 (jim_thompson5910):

`twice g` means `2 times g`

jimthompson5910 (jim_thompson5910):

so instead of g^2 you should have 2g

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

g^2 means "g squared" g^2 = g*g

OpenStudy (anonymous):

5h+2g=23?

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok....can u help with a couple more?

jimthompson5910 (jim_thompson5910):

I'll help with one more

OpenStudy (anonymous):

One forth the sum of r and ten is identical to r minus 4

OpenStudy (anonymous):

would you use a fraction for one forth...or literally use 1/4?

jimthompson5910 (jim_thompson5910):

1/4 is a fraction, so yes

jimthompson5910 (jim_thompson5910):

" the sum of r and ten" what does that translate to?

OpenStudy (anonymous):

+++++

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

Addition?

jimthompson5910 (jim_thompson5910):

yes so " the sum of r and ten" means r+10, agreed?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so 1/4 (x or +) r+10

jimthompson5910 (jim_thompson5910):

so 1/4 of that is simply \[\Large \frac{1}{4}(r+10)\]

jimthompson5910 (jim_thompson5910):

1/4 times (r+10) or 1/4 * (r+10)

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

\[\frac{ 1 }{ 4} \times r+10=r-4\]

OpenStudy (anonymous):

correct?

jimthompson5910 (jim_thompson5910):

don't forget the parenthesis around the "r+10"

OpenStudy (anonymous):

ohh ok

jimthompson5910 (jim_thompson5910):

the 1/4 is being multiplied by ALL of "r+10" and not just the r

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1 more?

OpenStudy (anonymous):

Degrees Kelvin K equals 273 plus degrees Celsius C

jimthompson5910 (jim_thompson5910):

what do you have for this one

OpenStudy (anonymous):

K+C=273

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

you have the equal sign in the wrong spot

jimthompson5910 (jim_thompson5910):

have a closer look at the original sentence

OpenStudy (anonymous):

k=273+c

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

the total cost of C is the price per gallon times the number of gallons g.

OpenStudy (anonymous):

GP=C